# thanks: Corresponding author

### v2

@@ V.2 Generalization across system sizes
A possible way to perform the training for this task would be to equalize the spin-opstring lengths of the larger and smaller systems through padding. However, this method is not compatible with arbitrary length data during inference and it also has an additional computational overhead. Instead, to handle the variable length data, we have incorporated the Global Pooling Layer (GPL) in 1D CNN architecture. In particular for the 1D SSH model we keep the same network structure as before but replace the flattening layer with a global average pooling layer. This layer reduces output of each CNN filter to a single value, which is an average over all its entries, thus effectively converting inputs of varying lengths into fixed-size output vectors. The subsequent FC layers always receive inputs of consistent size, allowing the model to generalize across different system sizes without needing padding or truncation.
In Fig. 11 we show the prediction of the ML model when trained on the data generated for a $24$-site periodic SSH chain and tested on a system with $32$ sites. We can see that the model almost accurately predicts the transition point with less than $3\%$ deviation. This demonstrates that our network is not rigidly tied to the specific length of the training data, but rather extracts scale-invariant patterns of the phases.
@@ VI Spin-state or opstring ?
Figure 12: Variance $\lambda_{1}$ of the first principal component obtained from PCA done on data collected from deep within the empty phase ($\rho=0$) and Mott insulator (MI) phase. The top plot shows the result of the analysis on the combined spin-opstring data while the bottom plot shows the case with only opstring data.
We have seen above that the combination of initial spin-state and opstring obtained from QMC simulations can be successfully used for classification, regression and transfer learning tasks. It would be interesting to ask whether only the spinstate or opstring may be used without the other for training and prediction. The operator string combined with a spin configuration of the lattice is equivalent to an entire QMC simulation cell which embodies quantum features unique to a particular phase. However, we find that even when trained solely on the opstring data a ML model is capable of predicting the phase transitions of the $2D$ $t_{2}-V_{1}$ model, discussed above in Sec. IV.1. But, the opstring may not be effective as a standalone training input between two phases which are related by some symmetry. For example, in the $2D$ $t_{2}-V_{1}$ model, the empty phase at $\rho=0$ and the Mott insulator at $\rho=1$, which are related by a particle-hole symmetry transformation, have similar QMC operator strings but very different spin-states. To illustrate this, in Fig. 12 we show the variance along first principal component from a principal component analysis (PCA) performed separately on spin-opstring and opstring data, collected deep within the empty and the MI phases. While the distinction between the combined spin-opstring in the two phases is evident, the operator strings exhibit similar features in both. This suggests that the opstring data alone may not always be sufficient for successful ML training.
Figure 13: Variance $\lambda_{1}$ along the first principal component obtained from PCA done on spin-opstring (Left) and spinstate (Right) data independently for values of coupling across the transition in the $1D$ SSH model.
It is well-known by now that spin configurations can be used to predict thermal phase transitions with great accuracy in classical spin systems. However, for quantum phase transitions, additional information beyond the spin-state may be necessary. As an example, let us consider the spin-state and spin-opstring data in the 1D SSH model. We show the results of a PCA performed on the spin-state data in Fig. 13 right and on the spin-opstring data in Fig. 13 left. The principal components (the variance of the first one is shown in Fig. 13) for the spin-state data fail to reveal any distinct features for a wide range of couplings across the transition, whereas the spin-opstring data exhibits a clear trend from which even the phase transition can be determined (by the location of the trough). This indicates that the spin-opstring data carries information regarding the transition in such a manner that can be detected by linear dimensionality reduction techniques like PCA.
Figure 14: Training and validation loss versus the number of epochs for a ML model trained with opstring data from the WTI model.
However, non-linear methods like neural networks might still be able to extract necessary information from the spin-state or opstring alone. To illustrate the advantage of the combined spin-opstring data using neural networks, we consider another system of hardcore bosons on a 2D periodic honeycomb lattice governed by the Hamiltonian,
$H=-t_{1}\sum_{\langle i,j\rangle}(b_{i}^{\dagger}b_{j}+\mathrm{h.c.})+\sum_{i}W_{i}n_{i}-\mu\sum_{i}n_{i}.$ (14)
@@
To demonstrate the advantage of spin-opstring data, we focus on the line at $W_{0}=0.5$ in the phase diagram, where as a function of the chemical potential $\mu$, the system goes through two phase transitions, from an empty phase at density $\rho=0$ to a superfluid phase, followed by a transition to a MI phase at density $\rho=1$. We have collected the training data from deep within the three phases and trained a ML model separately with three different input data types : opstring, spin-state and spin-opstring. The architecture of the ML model is same as the one used for predicting the phase boundaries of the $t_{2}-V_{1}$ model.
In Fig. 14, we show the training and validation losses as a function of the number of epochs for the ML model trained with only the opstring data. While the training loss decreases to $\mathcal{O}(10^{-5})$, we observe that the validation loss keeps growing, which is a classic sign of overfitting. The reason behind this can be understood in the following way. As the empty and MI phases are related by particle-hole symmetry, the opstring for these two phases are very similar. Therefore, when trained with these almost identical opstring data with different phase labels, the ML model attempts to minimize the loss on the training set by memorizing the contradictory labels. This prevents it from learning generalizable patterns, causing its performance to degrade on unseen validation data.
@@
The above result is, in fact, consistent with the conclusions in Ref. [40]. There it was observed that a ML model trained on density-density correlation function, which is a static quantity obtained from the spinstate, was unable to correctly predict the above transitions. On the other hand, the spatial correlation function, which encodes information of the temporal direction of the simulation cell, performed comparatively better [^2] It should be noted that the data for spatial correlation function in the empty and MI phases are related by particle-hole symmetry, similar to the opstring. Therefore, the prediction for the empty-SF and SF-MI transitions were performed separately with two separately trained ML models.. When trained with spin-opstring data, the ML model can, however, capture the transition points accurately as can be seen in Fig. 15. The combination of the spinstate and opstring therefore, proves effective in both avoiding data duplication and encoding quantum information necessary for the classification task. This brings us to the conclusion that while spin state or opstring data alone may suffice in some cases, the spin-opstring input proves to be more reliable and broadly applicable across different scenarios.
@@ VII Interpretability
In this work, we first employ Layer-wise Relevance Propagation (LRP), a powerful post hoc interpretability technique that decomposes the prediction of the model by propagating relevance backward from each output neuron to the input features [29]. With the application of a set of conservation-based propagation rules at each layer, LRP yields a relevance map over the input data.
$R_{j}=\sum_{k}\frac{z_{jk}}{\epsilon+\sum_{j}z_{jk}}R_{k}\,$, (15)
where $R_{j}$ corresponds to the score for a neuron $j$ in the inner layer and $R_{k}$ for a neuron $k$ in the outer layer. The coefficients $z_{jk}$ represent the model weights and biases present in between the two layers. The core idea of the LRP-$\epsilon$ rule involves adding a small positive constant $\epsilon$ to the denominator, the role of which is to offset some relevance from weak or contradictory neuron activations. Thus it helps to retain the most salient relevances by suppressing noise and avoids numerical instabilities arising from vanishing activations. For our LRP analysis, we have utilised the iNNvestigate library [41] which allows integration with Keras.
We start our interpretability analysis with the phase classification tasks for the 2D $t_{2}-V_{1}$ model. The trained ML model used for the prediction of the SF-cS phase transition (cf. Fig. 3) is analysed on the test dataset using the LRP-$\epsilon$ rule with $\epsilon=10^{-3}$. For better interpretability, the ML model without the final softmax activation layer is utilised for this purpose and we consider results with respect to the output neuron having the highest logit value. In Fig. 16 we show the raw relevance maps obtained for test data points within the SF and cS phases. Unlike relevance map visualizations of images like handwritten digits, the inherent complexity of the spin-opstring data makes direct interpretation challenging.
To extract physically meaningful insights, we utilize the underlying structure of the spin-opstring data, which is composed of two primary parts: the spin-state and the operator string. The operator string further comprises of three operator types, identity, diagonal, and off-diagonal. Based on this structure, we compute separate sums of the LRP scores for the spin-state, the operator string and each of the three individual operator types.
The total LRP sums for the spin-state and the operator string as a function of NN repulsion $V_{1}$, depicted in Fig. 17 top, exhibits a sharp jump precisely at the predicted transition point. Throughout this section, the error bars represent the standard deviation of a single model’s interpretability scores across the test samples at each coupling value. Quantitatively, the operator string provides the main contribution to the overall LRP, highlighting its dominant role in the SF-cS phase classification. For the SF phase, the spin-state gets zero relevance, which indicates that the machine relies solely on the operator string, due to the absence of any order in the spin-state. On the other hand in the cS phase the spin-state has a definite structure and the positive and negative relevances correspond to the two degenerate spin configurations.
To further understand the contributions of the different operator types, we plot the corresponding LRP sums in Fig. 17 bottom. Identity operators show zero relevance across the transition and are thus excluded from the figure. In contrast, both diagonal and off-diagonal operators exhibit clear jumps at the transition point. Diagonal operators are relevant in both the SF and cS phases, while off-diagonal operators contribute only within the SF phase, consistent with the physical nature of the two phases. Since superfluidity arises from particle hopping, which is absent in the insulating cS phase, off-diagonal operators naturally play a role only in the SF phase.
Figure 18: LRP analysis of the cS-cSS-MI transition as a function of chemical potential $\mu$ with repulsion fixed at $V_{1}=4.0$. Top: LRP sums for the spin-state and the operator string. Bottom: LRP contributions from diagonal and off-diagonal operators; identity operators are omitted due to zero relevance. The spin-state and off-diagonal sums are scaled for better visibility. In both plots, the predicted transition points are indicated by dashed vertical lines.
Next, we consider the cS-cSS-MI phase transition as a function of chemical potential $\mu$ (cf. Fig. 4). Similar to the SF-cS transition, the operator string remains the primary contributor in this case, as depicted in Fig. 18 top. In the plot for individual operator contributions (Fig. 18 bottom), identity operators are excluded due to zero relevance, and off-diagonal score sums have been rescaled for clarity. As expected from their insulating nature, the cS and MI phases show no relevance for off-diagonal operators. In the cSS phase, however, which is a homogeneous coexistence of superfluidity and cS, both the diagonal and off-diagonal operators yield finite relevance. The diagonal operators are also relevant in the MI phase and the increasing relevance sums with $\mu$ are actually proportional to the number of diagonal operators present in the opstring.
@@
However, in the cS phase, diagonal operators receive negative relevance, indicating that they provide evidence against the predicted phase. Interestingly, we also observe a negative total relevance sum in this phase which should have been positive and roughly equal to the logit value of the corresponding output neuron. The source of this missing relevance is unclear. To explore this issue further, we employed another widely used interpretability technique called SHapley Additive exPlanations (SHAP).
@@
SHAP is a model-independent method, based on cooperative game theory, where each input feature is treated as a “player” in a game [30]. The goal of this method is to fairly distribute the model’s prediction among the features according to their marginal contributions. This is done by measuring how the model’s output changes when the input features are systematically added or removed. In its general form (e.g., KernelSHAP), SHAP is model-agnostic, which makes it particularly valuable for validating interpretability results across a wide range of models. In our study, we utilize the Deep SHAP algorithm, a variant tailored for deep neural networks, to compute the SHAP scores, using the Python library based on [30].
The SHAP analysis for the cS-cSS-MI transition, shown in Fig. 19 top, reveals that the contribution of the spin-state is much smaller than the operator string, similar to the LRP analysis. One crucial difference between the LRP and SHAP analysis is the positive SHAP score for both spin-state and operator string in the cS phase, the sum of which is consistent with the logit value. In stark contrast to the LRP analysis, the identity operators contribute the most in the identification of the cS phase (see Fig. 19 bottom). This apparently resolves the ambiguity left by LRP, regarding which feature supports the prediction of the cS phase. For the cSS and MI phases, SHAP and LRP results are consistent.
Figure 20: SHAP score sums of individual operators for the SF–cS transition, as a function of increasing repulsion $V_{1}$. Predicted transition point is shown as the dashed vertical line.
@@
The contrasting relevance assigned to the identity operators by LRP and SHAP originates from the fundamental differences in how the two methods attribute relevance. In the spin-opstring data, the identity operators are encoded as zero. Since LRP propagates relevance based on neuron activations and weights, it tends to assign zero relevance to features that are numerically zero in the input [42], regardless of their contextual role. In contrast, SHAP evaluates the contribution of each feature by systematically perturbing the input and measuring its impact on the model’s output. This allows it to assign non-zero relevance to identity operators if their absence or presence influences the prediction.
Since identity operators are an integral part of our spin-opstring data, SHAP appears to be more suitable than LRP for our analysis. Thus, we use SHAP for the remainder of our interpretability studies. We re-analysed the predictions for the SF–cS transition using SHAP to investigate whether it offers additional insights beyond LRP. While both methods produce total relevance scores that closely match the model’s logit output, SHAP redistributes the importance across all three operators, including identity operators, as can be seen from Fig. 20. According to the SHAP analysis, in the cS phase, identity operators emerge as the dominant contributors to the model’s prediction. Insterestingly, this pattern is consistent across both the SF-cS and cS-cSS-MI transitions. This provides a coherent and physically meaningful picture of how the model identifies a phase.
Next, we performed SHAP analysis for the regression task in the cS-cSS-MI phase transitions with superfluid density $\rho_{s}$ as the regression target (cf. Fig. 5). Since the contribution from the spin-state is significantly smaller than the operator string, we focus on the individual contributions of different operator types in Fig. 21. In the cSS phase a pattern consistent with the classification task emerges : all operator types receive positive score. Since the cSS phase is associated with a finite $\rho_{s}$ value, it is expected that features indicative of this phase naturally contribute positively to a non-zero $\rho_{s}$ prediction. In contrast for the cS and MI phases, the SHAP results diverge from those seen in classification : identity operators now show negative SHAP values in the cS phase, while diagonal operators exhibit negative score in the MI phase. This discrepancy arises from the fundamentally different objectives of classification and regression. Classification predicts phase probabilities, while regression estimates a continuous value, in this case, the superfluid density. Since the cS and MI phases both exhibit zero $\rho_{s}$, features characteristic of these phases act to push the predicted $\rho_{s}$ value towards zero. Given that the average baseline prediction (influenced by the cSS phase) is positive, features that reduce the predicted value below this baseline naturally receive negative SHAP values. Thus, the negative scores of identity and diagonal operators in the cS and MI phase, respectively, do not imply irrelevance, but rather reflect their effectiveness in suppressing the predicted $\rho_{s}$ to zero.
@@
In Sec. V.1 we demonstrated the performance of the spin-opstring in a transfer learning setting, where a ML model was trained on labelled data from the non-interacting SSH model and evaluated on its interacting version. To understand what features the machine learns in both regimes, we have applied SHAP to two cases, predictions made on the non-interacting model, and on the interacting model with NN repulsion $V_{1}=1.0$. In Fig. 22 top, the SHAP value sums show that the spin-state and opstring contribute almost equally to the predictions in the non-interacting case. Fig. 22 middle further breaks down opstring contributions by operator type, revealing that diagonal operators carry zero SHAP value, while identity and off-diagonal operators display an inverse relevance pattern. Specifically, when one is positively relevant for a phase, the other consistently shows low or negative relevance, and vice-versa for the alternative phase.
@@
These observations offer a crucial insight into how the model distinguishes the phases of the SSH model. The diagonal operators encode static properties of the system which do not vary significantly across the transition in the non-interacting SSH model. Therefore, the ML model does not rely on them to identify either of the two phases, topological or trivial. On the other hand, off-diagonal operators represent hopping terms between sites, whereas identity operators signify the absence of hopping. The complementary behavior of these two operator types in the SHAP analysis indicates that the machine has learned the SSH model’s dimerization pattern, i.e., the alternating strong and weak bonds. For example, if a phase is characterized by strong hopping on certain bonds (e.g. the even bonds), the model assigns high importance to the off-diagonal operators. But if the machine encounters identity operators at those same locations, it makes the model less confident in that phase, resulting in negative relevance for the identity operators. The opposite pattern holds for the other phase. This clear connection between feature importance and the underlying physical phenomenon shows that the model is not merely fitting the data, it is learning the dimerization pattern that distinguishes the topological and trivial phases of the SSH model.
@@
Next, we observe that introducing interactions in the SSH model primarily affects the relevance of the diagonal operators, while the importance of other operator types remains largely unchanged. As shown in Fig. 22 bottom, the summed SHAP values for the interacting case ($V_{1}=1.0$) reveal a significant difference, i.e., the diagonal operators now have nonzero relevance, positive in one phase and negative in the other. This is an important finding which shows that the ML model, with its trained knowledge of the non-interacting system, has been able to utilise features in the interacting Hamiltonian. Even though the model never interpreted diagonal operators as meaningful during training, it correctly identifies their role during inference on interacting data, which is a hallmark of effective generalization in transfer learning. This behavior underscores that spin-opstring encodes the underlying physics in a way that allows a model to adapt.
@@
A recent interesting work [28] employed another interpretability technique, class activation mapping, to gain insights into CNN models trained on the SSH model used to predict the phase boundaries in the presence of disorder. The study aims to understand the poor generalization of CNN models in a transfer learning task, which was previously observed in [38]. Our study offers a complementary perspective in the interacting SSH model with a different input data, the spin-opstring, which allows succesful interpretability using simple ML models.

### v3

@@ V.2 Generalization across system sizes
For this purpose, to handle the variable length data, a Global Pooling Layer was incorporated into our 1D CNN architecture (further details can be found in Appendix C.) In Fig. 10 we show the prediction of the ML model when trained on the data generated for a $24$-site periodic SSH chain and tested on a system with $32$ sites. We can see that the model almost accurately predicts the transition point with less than $3\%$ deviation. This demonstrates that our network is not rigidly tied to the specific length of the training data, but rather extracts scale-invariant patterns of the phases.
@@ VI Spin-state or opstring ?
Figure 11: Variance $\lambda_{1}$ along the first principal component obtained from PCA done on spin-opstring (Left) and spin-state (Right) data independently for values of coupling across the transition in the $1D$ SSH model.
The operator string combined with the initial spin configuration, equivalent to an entire QMC simulation cell, embodies quantum features unique to a particular phase. We have seen above that this data representation obtained from QMC simulations can be successfully used for classification, regression and transfer learning tasks. However, it has been demonstrated in previous studies that the spin configuration alone can prove successful for the above ML tasks in various quantum systems. Therefore, it is important to ask whether there exist scenarios where the additional quantum information encoded in the spin-opstring becomes crucial for accurate phase recognition. In this section, we provide an example of a 2D HCB system where the spin-state alone turns out to be insufficient.
Before proceeding with the example, let us consider the spin-state and spin-opstring data in the already discussed 1D SSH model. We show the results of a principal component analysis (PCA) performed on the spin-state data in Fig. 11 right and on the spin-opstring data in Fig. 11 left. The principal components (the variance of the first one is shown in Fig. 11) for the spin-state data fail to reveal any distinct features for a wide range of couplings across the transition, whereas the spin-opstring data exhibits a clear trend from which even the phase transition can be determined (by the location of the trough). This indicates that spin-opstring carries information regarding the transition already in such a manner that can be detected by linear dimensionality reduction techniques like PCA.
However, non-linear methods like neural networks might still be able to extract necessary information from the spin-state. To illustrate the advantage of the combined spin-opstring data over the spin-state alone, using neural networks, we consider a system of HCBs on a 2D periodic honeycomb lattice governed by the Hamiltonian,
$H=-t_{1}\sum_{\langle i,j\rangle}(b_{i}^{\dagger}b_{j}+\textrm{h.c.})+\sum_{i}W_{i}n_{i}-\mu\sum_{i}n_{i}.$ (13)
@@
We focus on the line at $W_{0}=0.5$ in the phase diagram, where as a function of the chemical potential $\mu$, the system goes through two phase transitions, from an empty phase at density $\rho=0$ to a superfluid phase, followed by a transition to a MI phase at density $\rho=1$. We have collected the training data from deep within the three phases and trained a ML model separately with spin-state and spin-opstring as input data types. The architecture of the ML model is same as the one used for predicting the phase boundaries of the $t_{2}-V_{1}$ model.
@@
Figure 13: Variance $\lambda_{1}$ of the first principal component obtained from PCA done on data collected from deep within the empty phase and Mott insulator (MI) phase. The top plot shows the result of the analysis on the combined spin-opstring data while the bottom plot shows the case with only opstring data.
@@
The above result is, in fact, consistent with the conclusions in Ref. [41]. There it was observed that a ML model trained on density-density correlation function, which is a static quantity obtained from the spin-state, was unable to correctly predict the above transitions. On the other hand, the spatial correlation function, which encodes information of the temporal direction of the simulation cell, performed comparatively better [^2] It should be noted that the data for spatial correlation function in the empty and MI phases are related by particle-hole symmetry, similar to the opstring. Therefore, the prediction for the empty-SF and SF-MI transitions were performed separately with two separately trained ML models..
In contrast, when trained with spin-opstring data which encode both spatial and temporal information, the ML model captures the transition points accurately as can be seen in Fig. 12. We have rigorously tested the ML models across 100 independent training runs with random initializations. For every run, models trained on spin-opstring data consistently predict the phase transitions accurately, while those trained on spin-state fail to do so. Thus, our findings demonstrate the advantage of spin-opstring data over spin-state alone, establishing it as essential for accurate and reliable identification of phase transitions in general.
Furthermore, it would also be interesting to ask whether the opstring alone may suffice for ML tasks. In fact, we find that when trained solely on the opstring data, a ML model is capable of predicting the phase transitions of the 2D $t_{2}-V_{1}$ model, discussed above in Sec. IV.1. But, the opstring may not be effective as a standalone training input between two phases which are related by some symmetry. For example, in the 2D $t_{2}-V_{1}$ model, the empty phase at $\rho=0$ and the Mott insulator at $\rho=1$, which are related by a particle-hole symmetry transformation, have similar QMC operator strings but very different spin-states. To illustrate this, in Fig. 13 we show the variance along first principal component from PCA performed separately on spin-opstring and opstring data, collected deep within the empty and the MI phases. While the distinction between the combined spin-opstring in the two phases is evident, the operator strings exhibit similar features in both. This lack of distinction is further illustrated using neural networks in the following discussion.
In the 2D WTI model, we train the same ML model now with the opstring data alone to detect the phase transitions at $W_{0}=0.5$. During the training process, we observe that the training loss decreases to $\mathcal{O}(10^{-5})$, while the validation loss keeps growing, which is a classic sign of overfitting. The reason behind this can be understood in the following way. As the empty and MI phases are related by particle-hole symmetry, the opstring for these two phases are very similar. Therefore, when trained with these almost identical opstring data with different phase labels, the ML model attempts to minimize the loss on the training set by memorizing the contradictory labels. This prevents it from learning generalizable patterns, causing its performance to degrade on unseen validation data.
Thus, our study shows that the combination of the spinstate and opstring proves effective in both avoiding data duplication and encoding quantum information necessary for the classification task. This brings us to the conclusion that while spin state or opstring data alone may suffice in some cases, the spin-opstring input proves to be more reliable and broadly applicable across different scenarios.
@@ VII Interpretability
Figure 15: SHAP analysis for the SF-cS transition as a function of increasing repulsion $V_{1}$. Top: Score sums for spin-state (scaled for clarity) and operator string. Bottom: Score sums of individual operators. Predicted transition point is shown as the dashed vertical line.
@@
A key advantage of the spin-opstring representation is that its structured format naturally supports interpretability of ML predictions based on physical phenomena. The data consists of two components: the spin configuration and the operator string, which includes diagonal, off-diagonal, and identity operators, each associated with specific physical processes. This structure allows us to examine which features the model prioritizes in different phases. For example, in superfluid phase, the model focuses on off-diagonal operators that capture particle hopping, whereas in insulating phases, diagonal and identity operators become more relevant. Notably, in the 1D SSH model, where a local order parameter is absent, the network distinguishes the two dimerized phases through an inverse relevance pattern between off-diagonal and identity operators, which directly reflects the alternating bond structure central to SSH physics. Overall, the underlying structure of the spin-opstring enables meaningful insights into the model’s behavior, something not easily achieved with conventional inputs.
In this work, we employ two widely used interpretability techniques, Layer-wise Relevance Propagation (LRP) and SHapley Additive exPlanations (SHAP). Both these methods aim to explain a trained ML model’s prediction by distributing the final neuron outputs back to the input features, thereby quantifying each feature’s contribution. For example, in Fig. 14 we show the raw relevance or scores for input features using LRP, obtained for test data points within the SF and cS phases in the 2D $t_{2}-V_{1}$ model. Unlike relevance map visualizations of images like handwritten digits, a direct interpretation of individual features of the spin-opstring appears challenging. To extract physically meaningful insights, we leverage the underlying structure to compute separate sums of the LRP or SHAP scores for spin-state, opstring and each of the three individual operator types.
We observed a key difference in how each method attributes relevance. LRP, due to its reliance on input values for relevance propagation, assigns zero relevance to identity operators, which are encoded as zeros in the spin-opstring data. In contrast, SHAP can assign non-zero relevance to identity operators if their presence or absence influences the model’s output. Since identity operators are an integral part of our spin-opstring data, SHAP appears to be more suitable than LRP for our analysis. Therefore, in the main text, we present our interpretability results using SHAP, while details of the LRP-based analysis are included in Appendix D. Apart from the discrepancy regarding identity operators, the LRP results support the SHAP analysis presented below.
@@
SHAP is a model-independent method, based on cooperative game theory, where each input feature is treated as a “player” in a game [30]. It aims to fairly distribute the model’s prediction among the features by measuring how the model’s output changes when the input features are systematically added or removed. In its general form (e.g., KernelSHAP), SHAP is model-agnostic, which makes it particularly valuable for validating interpretability results across a wide range of models. In our study, we utilize the Deep SHAP algorithm, a variant tailored for deep neural networks, to compute the SHAP scores, using the Python library based on [30].
We start our interpretability analysis with the phase classification tasks for the 2D $t_{2}-V_{1}$ model. The trained ML models used for the prediction of various phase transitions is analysed on the corresponding test datasets. For better interpretability, the ML model without the final softmax activation layer is utilised for this purpose and we consider results with respect to the output neuron having the highest logit value. Throughout this section, the error bars represent the standard deviation of interpretability scores for a single model across multiple test samples at each coupling value.
The total SHAP scores for the spin-state and opstring, shown in Fig. 15 top as a function of $V_{1}$, reveal that the model primarily uses the opstring for SF-cS phase classification. Since the spin-state scores are nearly identical across both phases, it confirms that the model doesn’t rely on this feature. As depicted in Fig. 15 bottom, a closer look at the operator types shows that all of them exhibit clear jumps at the transition point. Diagonal operators are relevant in both phases, while off-diagonal operators contribute only in the SF phase, consistent with the presence of particle hopping. Within the insulating cS phase, identity operators become the dominant contributors to the model’s prediction.
The SHAP analysis for the cS-cSS-MI transition, reveals that as in the SF-cS case, the contribution of the spin-state is much smaller than the opstring (Fig. 16 top). Consistent with their insulating nature, the cS and MI phases show almost no relevance for off-diagonal operators as depicted in Fig. 16 bottom. In the cSS phase, however, which is a homogeneous coexistence of superfluidity and cS, both the diagonal and off-diagonal operators become important. Diagonal operators are also relevant in the MI phase and the increasing score sums with $\mu$ are actually proportional to the number of diagonal operators present in the opstring. Consistent with the cS-SF transition, identity operators contribute the most in the identification of the cS phase. This provides a coherent and physically meaningful picture of how the model identifies a phase.
@@
Next, we performed SHAP analysis for the regression task in the cS-cSS-MI phase transitions with superfluid density $\rho_{s}$ as the regression target (cf. Fig. 5). As the spin-state contribution is much smaller than the opstring, we focus on the individual operator types in Fig. 17. For the cSS phase, which has a finite $\rho_{s}$ value, all operator types contribute positively, which is consistent with the classification results (Fig. 16 bottom). However, for the cS and MI phases, SHAP assigns negative values for the key features identified in the classification task, i.e., identity operators for the cS phase and diagonal operators for the MI phase.
This discrepancy arises from the fundamentally different objectives of classification and regression. Since the cS and MI phases both exhibit zero $\rho_{s}$, features characteristic of these phases act to push the predicted $\rho_{s}$ value towards zero. Given a positive average prediction baseline set by the cSS data, features that reduce the predicted value below this baseline naturally receive negative SHAP values. Thus, the negative scores of identity and diagonal operators in the cS and MI phase, respectively, do not imply irrelevance, but rather reflect their effectiveness in suppressing the predicted $\rho_{s}$ to zero.
@@
In Sec. V.1 we demonstrated the performance of the spin-opstring in a transfer learning setting, where a ML model was trained on labelled data from the non-interacting SSH model and evaluated on its interacting version. To understand what features the machine learns in both regimes, we have applied SHAP to the non-interacting model as well as the interacting model (with $V_{1}=1.0$). In Fig. 18 top, the SHAP value sums show that the spin-state and opstring contribute almost equally to the predictions in the non-interacting case. Fig. 18 middle further breaks down opstring contributions by operator type, revealing that diagonal operators carry zero SHAP value, while identity and off-diagonal operators display an inverse relevance pattern. Specifically, when one is positively relevant for a phase, the other consistently shows low or negative relevance, and vice-versa for the alternate phase.
@@
These observations offer a crucial insight into how the ML model distinguishes the phases of the non-interacting SSH model. Diagonal operators encode static properties which do not vary significantly across the transition, so, the ML model does not rely on them to identify either of the two phases. On the other hand, off-diagonal operators represent hopping, while identity operators signify its absence. The complementary behavior of these two operators in the SHAP analysis indicates that the machine has learned the SSH model’s dimerization pattern, i.e., the alternating strong and weak bonds. For example, if a phase features strong hopping on even bonds, the model assigns high importance to the off-diagonal operators at those bonds. However, identity operators on even bonds makes the model less confident in that phase, resulting in their negative relevance. The opposite pattern holds for the other phase. This clear connection between feature importance and the underlying physical phenomenon shows that the model is not merely fitting the data, it is learning the dimerization pattern that distinguishes the topological and trivial phases of the SSH model.
@@
Introducing interactions in the SSH model primarily affects the relevance of diagonal operators, while the importance of other operators remains largely unchanged. In Fig. 18 bottom, the summed SHAP values for the interacting case ($V_{1}=1.0$) indeed reveal that the diagonal operators now have nonzero relevance, positive in one phase and negative in the other. This important finding shows that the ML model, with its trained knowledge of the non-interacting system, can utilise features in the interacting Hamiltonian. Although the model never interpreted diagonal operators as meaningful during training, it correctly identifies their role during inference on interacting data, a hallmark of effective generalization in transfer learning. This behavior underscores that spin-opstring encodes the underlying physics in a way that allows a model to adapt.
@@
A recent interesting work on the 1D disordered SSH model [Cybinski_2025] employed another interpretability technique, class activation mapping, to gain insights into trained CNN models. Our study offers a complementary perspective in the interacting SSH model with a different input data, the spin-opstring, which allows successful interpretability using simple ML models.